I need help with setting up this equation. A farmer has a total of 34 animals which comprise chickens and cows. Of all animals present, we counted 110 legs. How many of each animal does the farmer have? Please set up the equation and solve. My email address is hopealive@cox.net
Answer by sudhanshu_kmr(1152)


no. of chickens = x, now of cows = 34-x
no. of legs = 2*x + 4(34-x)
according to questi...
2x + 136 - 4x = 110
=> -2x = -26
=> x = 13
no. of chickens = 13, no. of cows = 21